OpenStudy (anonymous):

explain why 0/0=undefined.

OpenStudy (anonymous):

oh! no

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Leads to inconsistency...

OpenStudy (anonymous):

Prove 0=1 etc

OpenStudy (anonymous):

i can't understand

OpenStudy (anonymous):

if 0/0 is defined, 0*0/0=0*1/0. i.e: 0=1. satisfied , sritama?

OpenStudy (anonymous):

yes.thanks

OpenStudy (anonymous):

U can "talk about" 0/0 in analysis for limit purposes

OpenStudy (anonymous):

hey, e-studier, r u a teacher?

OpenStudy (anonymous):

I was, once....

OpenStudy (anonymous):

I am retired...

OpenStudy (anonymous):

professor?

OpenStudy (anonymous):

No, it was not my career..I did for a few years only....

OpenStudy (anonymous):

so, what do u do now? besides solving prolems, i mean

OpenStudy (anonymous):

Retirement is full-time,lol

OpenStudy (anonymous):

I study math, do 3D graphics, some other things...

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

is math your passion?

OpenStudy (anonymous):

sritama, day by day, u r becoming a philosopher

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

!!!

OpenStudy (anonymous):

u r putting some questions that can take even a scientist off guard

OpenStudy (anonymous):

Someone has to ask these questions, philosophy is cool...

OpenStudy (anonymous):

i think, philosophy is weird!

OpenStudy (anonymous):

oh . nice to think myself a philosopher!

OpenStudy (anonymous):

philosophy is good

OpenStudy (anonymous):

believe me, u r. if u keep it up, one day u will surpass karl marx :)

OpenStudy (anonymous):

thanks for ur imagination power dear.

OpenStudy (amistre64):

so why is 0/0 undefined; but 0! is defined, yet 0^0 is undefined ?? is my question. why do we define indeterminate forms for some aspects and not others?

OpenStudy (anonymous):

Convenience.

OpenStudy (amistre64):

thats what i thought :) \[\frac{d}{dx}(x) = 1x^0 \] is considered a special case becasue its tabboo to have x^0

OpenStudy (anonymous):

I didn't know that, I would have probably said it was 1 if asked.

OpenStudy (amistre64):

me too